OpenStudy (anonymous):

2+2=? @sam3d PLZ dont use calculator

OpenStudy (dmezzullo):

4

OpenStudy (anonymous):

4

hartnn (hartnn):

whats the point ?

OpenStudy (anonymous):

2+2=3

OpenStudy (dmezzullo):

lol

OpenStudy (anonymous):

I didnt close the question BUT IT IS CLOSED ........ how???

OpenStudy (anonymous):

2+2=? JOKE???

OpenStudy (dmezzullo):

lol good one

hartnn (hartnn):

some mod might have closed it, coz its not a serious question.....

OpenStudy (klimenkov):

2+2=11